OpenStudy (anonymous):

reduce the rational expression to its lowest terms 2x(2)+4x-126/3x(2)-243

OpenStudy (anonymous):

"2x(2)" is assumed to be twice x squared. \[\frac{2 x^2+4 x-126}{3 x^2-243}=\frac{(2 (-7+x) (9+x))}{(3 (-9+x) (9+x))}=\frac{2 (-7+x)}{3 (-9+x)} \]